OpenStudy (anonymous):

MEDAL GIVEN - Randall is saving for a $700 dirt bike. He has already saved $125 when he sees that the bike he wants is on sale for $550. He decides to ask his Dad for a loan to cover the rest of the cost. His father lends Randall the money at a simple interest rate of 3% per year. How much does Randall wind up paying to his Dad if he waits two years to pay back the entire loan plus interest? A. $437.75 B. $450.50 C. $566.50 D. $583.00

OpenStudy (anonymous):

@petrolinacarbon you know :o

OpenStudy (vanessagilbert1023):

B

OpenStudy (anonymous):

@vanessagilbert1023 can you explain

OpenStudy (vanessagilbert1023):

Randall borrows: 550-125 = $425 After 1 year the interest is 425 * 3/100 = $12.75 1st year is $425 + $12.75 = $437.75 Do same for second year Interest = $437.75 * 3/100 = $13.13 $437.75 + $13.13 = $450.88

OpenStudy (anonymous):

@vanessagilbert1023 thanks (: